If an object is thrown upward with an initial velocity of 96ft. per second, its height after 1 sec. is given by h=96t + 16t^2. Find the number of seconds before the object hits the ground.
Answer by Alan3354(69443)

If an object is thrown upward with an initial velocity of 96ft. per second, its height after 1 sec. is given by h=96t + 16t^2. Find the number of seconds before the object hits the ground.
-------------------
Your equation is not right, it's h(t) = 96t - 16t^2. If it's +16t^2, it would never stop going up.
h(t) = -16t^2 + 96t
It's a ground level when h(t) = 0
h(t) = -16t^2 + 96t = 0
16t^2 = 96t
t = 0 (at the start)
t = 6 seconds (at impact)